# Anatomical Foundation Models for Brain MRIs

### v3

@@ 3.3.3 Final objective function
The final objective function is a combination of our proposed AnatCL loss (either local or global) and the y-Aware loss [12] which considers age. Our final formulation is thus:
@@
where $\mathcal{L}_{age}$ is defined as in Eq. 1 and $\mathcal{L}_{AnatCL}$ is defined by either Eq. 3 or Eq. 5. The positive weights $\lambda_{1}$ and $\lambda_{2}$ control the contribution of each loss term independently. It is worth noting that the considered anatomical features in AnatCL can be computed directly from the 3D MRI with standard tools such as FreeSurfer [^1] https://surfer.nmr.mgh.harvard.edu/, thus our method does not require any additional label in the dataset.
@@ 4.1 Experimental data
We use a collection of T1-weighted MRI scans comprising 7,908 different individuals, for a total of 21,155 images. The experimental data is gathered from different publicly available datasets, and we target 5 different neurological conditions: healthy samples (OpenBHB [13]), Alzheimer’s Disease (ADNI [25] and OASIS-3 [22]), schizophrenia (SchizConnect [32]) and Autism Spectrum Disorder (ABIDE I [9]). The overall composition of the cohorts used in this study is presented in Fig. 2, while Tab. 1 provides an overview of all the downstream tasks that we explore in this work.
@@
OpenBHB OpenBHB is a recently released dataset that aggregates healthy control (HC) samples from many public cohorts (ABIDE 1, ABIDE 2, CoRR, GSP, IXI, Localizer, MPI-Leipzig, NAR, NPC, RBP). Every scan comes from a different subject, for a total of 3984 MRI images. We use OpenBHB as the pre-training dataset for our method. Besides the structural scans and patients information, OpenBHB also provides 7 anatomical measures according to the Desikan-Kiliani parcellation [7] (cortical thickness mean and std, gray matter volume, surface area, integrated mean, gaussian curvature index and intrinsic curvature index [13]).
@@
ADNI and OASIS-3 The Alzheimer’s Disease Neuroimaging Initiative (ADNI) is one of the largest and most popular datasets of MRIs for Alzheimer’s Disease (AD). In our study, we include all phases ADNI-1, ADNI-2, ADNI-GO and ADNI-3 amounting to 633 HC, 712 mild cognitive impairment (MCI) and 409 AD patients, for a total of 14,385 MRI images. We downloaded all raw acquisition in the DICOM format and performed VBM preprocessing with CAT12 [16]. We also collected all clinical information and processed it in order to distinguish between stable and progressive MCI. Additionally, we also employ the OASIS3 dataset, including 685 patients (88 AD cases and 597 control cases), for a total of 1354 additional images.
@@
SchizConnect For schizophrenia we employ the SchizConnect, a large dataset of MRIs. We include structural MRIs for a total of 383 patients, divided into 180 HC, 102 with schizophrenia (broad), 74 with schizophrenia (strict), 11 schizoaffective patients, and 9 with bipolar disorder, for a total of 383 MRIs.
@@
ABIDE I The Autism Brain Imaging Data Exchange (ABIDE) dataset is one of the most popular datasets of patients affected by conditions of Autism Spectrum Disorder (ASD). We leverage data from ABIDE-I, including structural MRIs for a total of 1102 patients, divided into 556 HC, 339 patients with autism, 93 patients with Asperger’s Syndrome, and 7 with pervasive developmental disorder not otherwise specified (PDD-NOS). The total number of images is 1102.
@@
Clinical assessment scores and phenotypes In addition to predicting the above diagnosis, we also consider some relevant clinical assessments included in the data, as reported in Tab. 1. For SchizConnect, we consider the Abnormal Involuntary Movement Scale (AIMS) evaluated in three assessments (overall, upper body, lower body), a Depression score based on the Calgary Scale, Handedness information, and GAIT measurements with the Simpson-Angus-Scale (SAS). For ABIDE, we also consider handedness and three IQ scores, namely Full Scale IQ (FIQ), Visual IQ (VIQ), and Performance IQ (FIQ) measured with the Wechsler Abbreviated Scale (WASI). To the best of our knowledge, this is the first work attempting to tackle this objective.
@@
4.1.1 Data preparation and preprocessing
@@
All the images underwent the same standard VBM preprocessing, using CAT12 [16], which includes non-linear registration to the MNI template and gray matter (GM) extraction. The final spatial resolution is 1.5mm isotropic and the images are of size 121 x 145 x 121. The preprocessing is performed using the brainprep package  [^2] https://brainprep.readthedocs.io/en/latest/. After preprocessing, we consider modulated GM images, as in [13].
4.1.2 Computation of anatomical features
Anatomical measurements on OpenBHB are released as part of the original dataset. All details can be found in [13]. For deriving the measuremets, the brainprep [^3] https://github.com/neurospin-deepinsight/brainprep module was employed, with FreeSurfer version FSL-6.0.5.1, and Desikan-Killiany atlas as provided in FreeSurfer [^4] https://surfer.nmr.mgh.harvard.edu/fswiki/CorticalParcellation.
@@ 4.1.2 Computation of anatomical features
4.2 Experimental setup
Our experiments involved pre-training a model using the proposed AnatCL loss on the OpenBHB dataset, followed by testing on various downstream tasks across different datasets. The experimental settings for the two loss formulations, local (AnatCL-G3) and global (AnatCL-L3), are identical. We pre-train two ResNet-18 3D models using VBM-preprocessed images and their corresponding Desikan measures with the proposed formulations. The training process employs the Adam optimizer with a learning rate of 0.0001, and a decay rate of 0.9 applied after every 10 epochs. The models are trained with a batch size of 32 for a total of 300 epochs. As values of $\lambda_{1}$ and $\lambda_{2}$, for simplicity, we use 1. As standard practice in CL approaches [5, 21], the contrastive loss is computed on a fully-connected projection head following the encoder, composed of two layers. To ensure robust evaluation, we perform cross-validation using 5 folds. The results are computed in terms of mean and standard deviation across the 5 folds. After the pre-training step we evaluate the models by testing their performance with a transfer learning approach: we extract the latent representations generated by the model using only the encoder of the model (i.e., discarding the fully connected head). For each downstream task, we train different linear classifiers on the extracted representations to assess the model’s ability to learn meaningful and generalizable features. To run our experiments, we employ a cluster of 4 NVIDIA V100 GPUs, with a single training taking around 10h.
4.2.1 Choice of Anatomical Features
OpenBHB provides 7 different anatomical measures, computed with FreeSurfer: Cortical Thickness (CT, mean and std values), Gray Matter Volume (GMV), Surface Area, Integrated Mean, Gaussian Curvature Index, and Intrinsic Curvature Index. To motivate the choice of the anatomical features, we conduct a preliminary study by training a simple linear model (Ridge regression) over each feature, with the task of brain age prediction on the OpenBHB data. The results are evaluated in terms of negative mean absolute error and $R^{2}$ score and are reported in Tab. 2. The best results are achieved by GVM and Surface Area, which are morphological features regarding the volume and the total area of each ROI. Thus, we select these two features for our experiments with AnatCL. In addition to GMV and Surface Area measures, we also include CT as it is a widely used measure in related works [29, 30]. The final choice is thus composed of three features, which cover different aspects of the morphology and geometry of brain ROIs (e.g. cortical thickness is relevant for conditions such as Alzheimer’s Disease [26], and volume and surface measurements can signal brain atrophy, linked to different conditions).
4.2.2 Baselines for comparison
In this work we focus on contrastive learning has it has been demonstrated to work better than non-contrastive method in recent literature [14]. For this reason, we pick recent state-of-the-art contrastive methods in neuroimaging, namely y-Aware [12] and ExpW [1]. In addition to these recent techniques, we also consider the popular SimCLR framework [5] as self-supervised contrastive baseline. For comparison with traditional non-contrastive supervised techniques, we also compare to baseline models trained with L1 loss for brain age prediction [20].
4.3 Brain age prediction
The first two downstream tasks we evaluate are brain age prediction and sex classification, on the OpenBHB dataset. The results are reported in Tab. 3. From the results, we conclude that AnatCL is able to slightly surpass state-of-the-art performance on brain age prediction, reaching the lowest absolute error of 2.61 years. It is worth noting that we do not employ any bias-correction method as in [6, 23], but we consider the raw predictions of the model. For sex classification, we do not match ExpW [1], but we improve over the SimCLR and L1 baselines.
Method Age MAE Sex
SimCLR [5] 5.58$\pm$0.53 76.7$\pm$1.67
L1 (age supervised) 2.73$\pm$0.14 76.7$\pm$0.67
y-Aware [12] 2.66$\pm$0.06 79.6$\pm$1.13
ExpW [1] 2.70$\pm$0.06 80.3$\pm$1.7
AnatCL-G3 2.61$\pm$0.08 78.2$\pm$1.25
AnatCL-L3 2.64$\pm$0.07 78.2$\pm$0.7
Table 3: Results on OpenBHB in terms of mean absolute error (MAE) on age prediction, and balanced accuracy on sex classification.
Schizconnect ABIDE-I
@@ 4.3 Brain age prediction
Method SCZ. (Broad) SCZ (Strict) Schizoaff. Bipolar Autism Aspergers PDD-NOS
SimCLR [5] 58.53$\pm$3.52 68.47$\pm$8.47 51.23$\pm$11.94 67.34$\pm$9.96 54.45$\pm$2.99 59.61$\pm$2.72 52.12$\pm$6.62
L1 65.79$\pm$4.74 70.68$\pm$5.10 65.24$\pm$15.21 64.49$\pm$22.08 54.53$\pm$1.79 61.33$\pm$8.77 57.54$\pm$5.93
y-Aware [12] 66.20$\pm$4.50 71.04$\pm$2.31 70.29$\pm$14.73 63.95$\pm$19.81 55.84$\pm$3.37 60.60$\pm$9.22 56.17$\pm$10.17
ExpW [1] 69.53$\pm$4.43 67.65$\pm$8.27 63.26$\pm$18.06 74.34$\pm$18.95 58.50$\pm$2.41 58.68$\pm$3.82 58.31$\pm$4.33
AnatCL-G3 72.55$\pm$5.16 71.03$\pm$8.53 73.65$\pm$7.29 63.43$\pm$15.94 53.48$\pm$0.99 59.32$\pm$6.58 53.38$\pm$5.86
AnatCL-L3 66.38$\pm$5.96 72.07$\pm$8.42 68.37$\pm$8.60 56.60$\pm$11.48 55.85$\pm$1.02 60.40$\pm$2.29 65.37$\pm$6.73
Table 4: Results on schizophrenia detection (SCZ) and autism spectrum disorder (ASD) in terms of balanced accuracy.
4.4 Schizophrenia and ASD
We evaluate downstream performance on SchizConnect for detecting schizophrenia (broad and strict), schizoaffective and bipolar patients. Results are reported in Tab. 4. With AnatCL we achieve state-of-the-art performance on three out of four tasks, denoting that taking anatomical information into account can prove useful for these psychiatric conditions. We also assess detection performance for Autism Spectrum Disorder (ASD) patients, across three categories (autism, aspergers and PDD-NOS), results are reported in Tab. 4. While AnatCL does not beat the other methods on autistic and aspergers patients, we significantly improve the accuracy for PDD-NOS patients, which is a rarer diagnosis in the data. Overall, AnatCL is able to more consistently achieve improved results than any other baseline.
4.5 Alzheimer’s desease and cognitive impariments
ADNI OASIS-3
Method HC vs AD sMCI vs pMCI HC vs AD
SimCLR [5] 78.47$\pm$2.51 61.77$\pm$3.85 73.97$\pm$4.98
L1 (age supervised) 81.20$\pm$2.3 68.12$\pm$5.42 75.40$\pm$5.4
y-Aware [12] 80.3$\pm$1.8 64.72$\pm$4.43 76.70$\pm$3.30
ExpW [1] 81.84$\pm$2.95 66.54$\pm$5.64 74.67$\pm$2.87
AnatCL-G3 80.47$\pm$2.95 66.03$\pm$2.93 75.59$\pm$2.67
AnatCL-L3 80.11$\pm$1.0 62.83$\pm$4.5 75.88$\pm$3.0
Table 5: Results on Alzheimer’s Disease (AD) classification in terms of balanced accuracy.
In Tab. 5 we report results for Alzheimer’s Disease (AD) detection on ADNI and OASIS-3. While we do not reach the best results overall, AnatCL can consistenly improve over the self-supervised baseline SimCLR and over the weakly supervised method y-Aware and reaches overall competitive results. On OASIS-3, AnatCL shows also better results than ExpW, in both local and global variants. Overall, AnatCL provides more consistent performance than other methods.
4.6 Clinical assessments and phenotypes
(a) ABIDE-I
(b) SchizConnect
Figure 4: Assessment scores and phenotypes prediction from brain MRIs.
As final experiments, we turn our attention to predicting clinical assessments score from brain MRIs. To the best of our knowledge, this has not been explored in other works, and it could provide useful insights on the relation between brain anatomy and behavioral phenotypes. The 10 phenotypes considered can be distinguished based on the prediction task: AIMS, depression, handedness and GAIT are classification tasks, while IQ scores (FIQ, VIQ and PIQ) are regression tasks. For handendess, we predict right-handed vs other (left-handed or ambi), for depression we classify between absent vs mild and above, for AIMS we classify between none and minimal vs mild and above, for GAIT between normal vs everything else. For more detailed explanation of the possible value in the considered phenotypes we refer to the official documentation [^5] http://schizconnect.org/documentation#data_dictionaries [^6] http://fcon_1000.projects.nitrc.org/indi/abide/ABIDE_LEGEND_V1.02.pdf. The results are reported in Fig. 4. While we cannot conclude that any of the analysed method can accurately predict all clinical assessments from MRI scans, AnatCL overall achieves the best results three out of ten times, which is more than any other baseline. Interestingly, AnatCL can better predict the overall AIMS score and patients’ handedness, hinting that brain anatomy may be linked with these phenotypes. The link between brain anatomy and involuntary movements could lead to the discovery of novel biomarkers and should be studied more in depth in future research.

### v4

@@ 3.3.1 Final objective function
The final objective function is a combination of our proposed AnatCL loss (either local or global) and the y-Aware loss which considers age. Our final formulation is thus:
@@
where $\mathcal{L}_{age}$ is defined as in Eq. 1 and $\mathcal{L}_{AnatCL}$ is defined by either Eq. 3 or Eq. 5. The weights $\lambda_{1}$ and $\lambda_{2}$ control the contribution of each loss term independently. It is worth noting that the considered anatomical features in AnatCL can be computed directly from the MRI with standard tools such as FreeSurfer [^2] https://surfer.nmr.mgh.harvard.edu/, thus our method does not require any additional label in the dataset.
@@ 4.1 Experimental data
We use a collection of T1-weighted MRI scans comprising 7,908 different individuals, for a total of 21,155 images. The experimental data is gathered from different publicly available datasets, and we target 5 different neurological conditions: healthy samples (OpenBHB (Dufumier et al., 2022)), Alzheimer’s Disease (ADNI (Petersen et al., 2010) and OASIS-3 (LaMontagne et al., 2019)), schizophrenia (SchizConnect (Wang et al., 2016)) and Autism Spectrum Disorder (ABIDE I (Di Martino et al., )). Tab. 1 provides an overview of all the downstream tasks that we explore in this work.
OpenBHB
@@
OpenBHB is a recently released dataset that aggregates healthy control (HC) samples from many public cohorts (ABIDE 1, ABIDE 2, CoRR, GSP, IXI, Localizer, MPI-Leipzig, NAR, NPC, RBP). Every scan comes from a different subject, for a total of 3984 MRI images. We use OpenBHB as the pre-training dataset for our method. Besides the structural scans and patients information, OpenBHB also provides 7 anatomical measures: cortical thickness mean and std, gray matter volume, surface area, integrated mean, gaussian curvature index, and intrinsic curvature index (Dufumier et al., 2022).
ADNI and OASIS-3
@@
The Alzheimer’s Disease Neuroimaging Initiative (ADNI) [^3] Data used in the preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). is one of the largest and most popular datasets of MRIs for Alzheimer’s Disease (AD). In our study, we include all phases ADNI-1, ADNI-2, ADNI-GO and ADNI-3 amounting to 633 HC, 712 mild cognitive impairment (MCI) and 409 AD patients, for a total of 14,385 MRI images. We downloaded all raw acquisition in the DICOM format and performed VBM preprocessing with CAT12 (Gaser et al., 2022). We also collected all clinical information and processed it in order to distinguish between stable and progressive MCI. Additionally, we also employ the OASIS3 dataset, including 685 patients (88 AD cases and 597 control cases), for a total of 1354 images.
SchizConnect
@@
For schizophrenia, we employ SchizConnect, a large dataset of MRIs. We include structural MRIs for a total of 383 patients, divided into 180 HC, 102 with schizophrenia (broad), 74 with schizophrenia (strict), 11 schizoaffective patients, and 9 with bipolar disorder, for a total of 383 MRIs.
ABIDE I
@@
The Autism Brain Imaging Data Exchange (ABIDE) dataset is one of the most popular datasets of patients affected by conditions of Autism Spectrum Disorder (ASD). We leverage data from ABIDE-I, including structural MRIs for a total of 1102 patients, divided into 556 HC, 339 patients with autism, 93 patients with Asperger’s Syndrome, and 7 with pervasive developmental disorder not otherwise specified (PDD-NOS). The total number of images is 1102.
Clinical assessment scores and phenotypes
@@
In addition to predicting the diagnosis, we also consider some relevant clinical assessments included in the data, as reported in Tab. 1. For SchizConnect, we consider the Abnormal Involuntary Movement Scale (AIMS) evaluated in three assessments (overall, upper body, lower body), a Depression score based on the Calgary Scale, Handedness information, and GAIT measurements with the Simpson-Angus-Scale (SAS). For ABIDE, we also consider handedness and three IQ scores, namely Full Scale IQ (FIQ), Visual IQ (VIQ), and Performance IQ (FIQ) measured with the Wechsler Abbreviated Scale (WASI). To the best of our knowledge, this is the first work attempting to tackle this objective.
@@
Data preparation and preprocessing
@@
All the images underwent the same standard VBM preprocessing, using CAT12 (Gaser et al., 2022), which includes non-linear registration to the MNI template and gray matter (GM) extraction. The final spatial resolution is 1.5mm isotropic and the images are of size 121 x 145 x 121. The preprocessing is performed using the brainprep package [^4] https://brainprep.readthedocs.io/en/latest/. After preprocessing, we consider modulated GM images, as in Dufumier et al. (2022).
Computation of anatomical features
Anatomical measurements on OpenBHB are released as part of the original dataset. For deriving the measurements, the brainprep module was employed, with FreeSurfer version FSL-6.0.5.1, and the Desikan-Killiany and Destrieux atlases as provided in FreeSurfer. We chose to employ CT, GMV, and SA as they are also used in Sebenius et al. (2023); for additional results on other measures, please refer to the supplementary materials.
4.2 Experimental setup
Our experiments involved pre-training a model using the proposed AnatCL loss on the OpenBHB dataset, followed by testing on various downstream tasks across different datasets. The experimental settings for the two loss formulations, local (AnatCL-G3) and global (AnatCL-L3), are identical. We pre-train two ResNet-18 3D models using VBM-preprocessed images and their corresponding anatomical measures with the proposed formulations. The training process employs the Adam optimizer with a learning rate of 0.0001 and a decay rate of 0.9 applied after every 10 epochs. The models are trained with a batch size of 32 for a total of 300 epochs. As values of $\lambda_{1}$ and $\lambda_{2}$, for simplicity, we use 1. As standard practice in CL approaches (Chen et al., 2020; Khosla et al., 2020), the contrastive loss is computed on a fully-connected projection head following the encoder, composed of two layers. To ensure robust evaluation, we perform cross-validation using 5 folds. The results are computed in terms of mean and standard deviation across the 5 folds. After the pre-training step we evaluate the models by testing their performance with a transfer learning approach: we extract the latent representations generated by the model using only the encoder of the model (i.e., discarding the fully connected head). For each downstream task, we train different linear classifiers on the extracted representations to assess the model’s ability to learn meaningful and generalizable features. To run our experiments, we use the Jean Zay and Leonardo clusters (providing NVIDIA V100 and A100 GPUs), with a single training taking around 10h.
Baselines for comparison
In this work we focus on contrastive learning has it has been demonstrated to work better than non-contrastive method in recent literature (Dufumier et al., 2024). For this reason, we pick recent state-of-the-art contrastive methods in neuroimaging, namely y-Aware (Dufumier et al., 2021) and ExpW (Barbano et al., 2023). In addition to these recent techniques, we also consider the popular SimCLR framework (Chen et al., 2020) as a self-supervised contrastive baseline (Tak et al., 2024). For comparison with traditional non-contrastive supervised techniques, we also compare with baseline models trained with L1 loss for brain age prediction (Jónsson et al., 2019).
4.3 Results
We evaluate the results of AnatCL pre-trained on OpenBHB and tested on different datasets, for several deep phenotyping and clinical assessment scores prediction tasks.
Brain age and sex prediction
The preliminary downstream tasks we evaluate are brain age prediction and sex classification, on OpenBHB. The results are reported in Tab. 2. AnatCL achieves state-of-the-art results on brain age prediction, reaching the lowest absolute error of 2.55 years. It is worth noting that we do not employ any bias-correction method (de Lange & Cole, 2020), but we consider the raw predictions of the model. For sex classification, AnatCL is competitive with ExpW (Barbano et al., 2023), and improves over the other baselines.
Table 2: Results on OpenBHB in terms of mean absolute error (MAE) on age prediction, and balanced accuracy on sex classification. AnatCL-L3 refers to the local version, while AnatCL-G3 refers to the global version; “3” refers to the three measures employed (CT, GVM, SA).
Method Age MAE Sex
SimCLR Chen et al. (2020) 5.58$\pm$0.53 76.7$\pm$1.67
L1 (age supervised) 2.73$\pm$0.14 76.7$\pm$0.67
y-Aware Dufumier et al. (2021) 2.66$\pm$0.06 79.6$\pm$1.13
ExpW Barbano et al. (2023) 2.70$\pm$0.06 80.3$\pm$1.7
AnatCL-L3 (Desikan) 2.55$\pm$0.07 80.0$\pm$1.0
AnatCL-L3 (Destrieux) 2.58$\pm$0.08 80.0$\pm$1.0
AnatCL-G3 (Desikan) 2.56$\pm$0.06 79.0$\pm$1.0
AnatCL-G3 (Destrieux) 2.59$\pm$0.05 80.0$\pm$1.0
Table 3: Deep phenotyping results across psychiatric (SCZ, ASD) and neuro-degenerative (AD) benchmarks in terms of balanced accuracy.
Schizconnect ABIDE-I ADNI OASIS-3
Method SCZ (Broad) SCZ (Strict) Schizoaff. Bipolar Autism Aspergers PDD-NOS HC vs AD sMCI vs pMCI HC vs AD
SimCLR 58.53$\pm$3.52 68.47$\pm$8.47 51.23$\pm$11.94 67.34$\pm$9.96 54.45$\pm$2.99 59.61$\pm$2.72 52.12$\pm$6.62 78.47$\pm$2.51 61.77$\pm$3.85 73.97$\pm$4.98
L1 (age-sup.) 65.79$\pm$4.74 70.68$\pm$5.10 65.24$\pm$15.21 64.49$\pm$22.08 54.53$\pm$1.79 61.33$\pm$8.77 57.54$\pm$5.93 81.20$\pm$2.30 68.12$\pm$5.42 75.40$\pm$5.40
y-Aware 66.20$\pm$4.50 71.04$\pm$2.31 70.29$\pm$14.73 63.95$\pm$19.81 55.84$\pm$3.37 60.60$\pm$9.22 56.17$\pm$10.17 80.30$\pm$1.80 64.72$\pm$4.43 76.70$\pm$3.30
ExpW 69.53$\pm$4.43 67.65$\pm$8.27 63.26$\pm$18.06 74.34$\pm$18.95 58.50$\pm$2.41 58.68$\pm$3.82 58.31$\pm$4.33 81.84$\pm$2.95 66.54$\pm$5.64 74.67$\pm$2.87
AnatCL-L3 (Desikan) 66.03$\pm$4.06 69.86$\pm$11.20 52.72$\pm$6.64 63.12$\pm$21.50 53.99$\pm$2.60 61.44$\pm$4.33 56.67$\pm$7.43 79.71$\pm$2.51 64.78$\pm$3.57 74.26$\pm$3.32
AnatCL-L3 (Destrieux) 67.31$\pm$5.75 70.03$\pm$7.16 72.47$\pm$12.27 55.88$\pm$13.23 53.14$\pm$4.20 60.99$\pm$4.94 56.63$\pm$6.74 79.38$\pm$3.56 64.97$\pm$3.40 77.00$\pm$2.28
AnatCL-G3 (Desikan) 64.31$\pm$4.43 72.46$\pm$4.41 62.42$\pm$14.95 72.42$\pm$15.77 57.12$\pm$2.79 58.25$\pm$4.46 64.39$\pm$9.26 80.84$\pm$2.38 64.88$\pm$4.37 78.32$\pm$4.48
AnatCL-G3 (Destrieux) 64.12$\pm$3.18 73.04$\pm$4.42 59.76$\pm$15.20 84.19$\pm$7.10 55.41$\pm$4.55 58.96$\pm$7.52 58.10$\pm$4.63 80.75$\pm$3.35 65.91$\pm$2.23 78.66$\pm$2.89
Schizophrenia and ASD
We evaluate phenotyping performance on SchizConnect for detecting schizophrenia (broad and strict), schizoaffective, and bipolar patients. Results are reported in Tab. 3. With AnatCL, we achieve state-of-the-art performance on three out of four tasks, suggesting that including anatomical information during training can prove useful for these psychiatric conditions. We also assess detection performance for Autism Spectrum Disorder (ASD) patients, across three categories (autism, aspergers, and PDD-NOS). While AnatCL does not beat the other methods on general autism (but is competitive), we achieve the best results on Aspergers and significantly improve the accuracy for PDD-NOS patients, which is a rarer diagnosis in the data. Overall, AnatCL achieves better results than other baselines most of the time, and is competitive with SOTA results in the remaining instances.
Alzheimer’s desease and cognitive impariments
In Tab. 3 we also report results for Alzheimer’s Disease (AD) detection on both ADNI and OASIS-3. On ADNI, we also assess the discriminative power between stable MCI (sMCI) and progressive MCI (pMCI) patients. While we do not reach the best results overall on ADNI, AnatCL can consistently improve over the self-supervised baseline SimCLR and over the weakly supervised method y-Aware, and reaches overall competitive results. On OASIS-3, AnatCL also achieves the best results in AD detection, in both local and global variants. Overall, AnatCL provides more consistent performance than other methods.
Clinical assessments scores
(a) ABIDE-I
(b) SchizConnect
Figure 3: Prediction of clinical assessment scores from structural brain MRIs.
As final experiments, we turn our attention to predicting clinical assessment scores, cognition, and psychopathology from brain MRIs. To the best of our knowledge, this has not been explored in other works, and it could provide useful insights on the relation between brain anatomy and individual phenotypes, whereas most of the literature focuses on fMRI (Kong et al., 2023). The 10 scores considered can be distinguished based on the prediction task: AIMS, depression, handedness, and GAIT are classification tasks, while IQ scores (FIQ, VIQ, and PIQ) are regression tasks. For handendess, we predict right-handed vs other (left-handed or ambi), for depression, we classify between absent vs mild and above, for AIMS, we classify between none and minimal vs mild and above, for GAIT, between normal vs everything else. For a more detailed explanation of the considered phenotypes, we refer to the official documentation. The results are reported in Fig. 3. While we cannot conclude that any of the analysed methods can accurately predict all clinical assessments from structural MRI scans, AnatCL overall achieves the best results three out of ten times, which is more than any other baseline. Interestingly, AnatCL can better predict the overall AIMS score (Fig. 3(b)), hinting that it may be linked to brain anatomy. This link could lead to the discovery of novel biomarkers and should be studied more in depth in future research.
5 Conclusions
We propose AnatCL, a foundation model based on brain anatomy and trained with a weakly contrastive learning approach. With thorough validation on 10 different downstream tasks, we show that incorporating anatomy information during training can result in more accurate predictions of different neurological and psychiatric conditions, and appears promising for clinical assessment scores. We release the weights of the trained model for public use, enabling researchers and practitioners to leverage AnatCL in many different applications.
Acknowledgments
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy). This work was granted access to the HPC/AI resources of IDRIS under the allocation 2023-AD011013473R1 made by GENCI. Data collection and sharing for the Alzheimer’s Disease Neuroimaging Initiative (ADNI) is funded by the National Institute on Aging (National Institutes of Health Grant U19AG024904).
Supplementary Material
6 Limitations, impact, and ethics
7 Results
7.1 Choice of Anatomical Features
@@
Table 5: AnatCL results with all seven anatomical features.
Schizconnect ABIDE-I ADNI OASIS-3
@@
Method SCZ (Broad) SCZ (Strict) Schizoaff. Bipolar Autism Aspergers PDD-NOS HC vs AD sMCI vs pMCI HC vs AD
SimCLR 58.53$\pm$3.52 68.47$\pm$8.47 51.23$\pm$11.94 67.34$\pm$9.96 54.45$\pm$2.99 59.61$\pm$2.72 52.12$\pm$6.62 78.47$\pm$2.51 61.77$\pm$3.85 73.97$\pm$4.98
L1 (age-sup.) 65.79$\pm$4.74 70.68$\pm$5.10 65.24$\pm$15.21 64.49$\pm$22.08 54.53$\pm$1.79 61.33$\pm$8.77 57.54$\pm$5.93 81.20$\pm$2.30 68.12$\pm$5.42 75.40$\pm$5.40
y-Aware 66.20$\pm$4.50 71.04$\pm$2.31 70.29$\pm$14.73 63.95$\pm$19.81 55.84$\pm$3.37 60.60$\pm$9.22 56.17$\pm$10.17 80.30$\pm$1.80 64.72$\pm$4.43 76.70$\pm$3.30
ExpW 69.53$\pm$4.43 67.65$\pm$8.27 63.26$\pm$18.06 74.34$\pm$18.95 58.50$\pm$2.41 58.68$\pm$3.82 58.31$\pm$4.33 81.84$\pm$2.95 66.54$\pm$5.64 74.67$\pm$2.87
AnatCL-L3 (Desikan) 66.03$\pm$4.06 69.86$\pm$11.20 52.72$\pm$6.64 63.12$\pm$21.50 53.99$\pm$2.60 61.44$\pm$4.33 56.67$\pm$7.43 79.71$\pm$2.51 64.78$\pm$3.57 74.26$\pm$3.32
AnatCL-L3 (Destrieux) 67.31$\pm$5.75 70.03$\pm$7.16 72.47$\pm$12.27 55.88$\pm$13.23 53.14$\pm$4.20 60.99$\pm$4.94 56.63$\pm$6.74 79.38$\pm$3.56 64.97$\pm$3.40 77.00$\pm$2.28
AnatCL-G3 (Desikan) 64.31$\pm$4.43 72.46$\pm$4.41 62.42$\pm$14.95 72.42$\pm$15.77 57.12$\pm$2.79 58.25$\pm$4.46 64.39$\pm$9.26 80.84$\pm$2.38 64.88$\pm$4.37 78.32$\pm$4.48
AnatCL-G3 (Destrieux) 64.12$\pm$3.18 73.04$\pm$4.42 59.76$\pm$15.20 84.19$\pm$7.10 55.41$\pm$4.55 58.96$\pm$7.52 58.10$\pm$4.63 80.75$\pm$3.35 65.91$\pm$2.23 78.66$\pm$2.89
AnatCL-L7 (Desikan) 66.16$\pm$6.12 70.92$\pm$8.82 63.65$\pm$18.60 66.81$\pm$20.04 55.28$\pm$1.41 60.76$\pm$7.85 59.42$\pm$8.02 80.17$\pm$1.98 64.59$\pm$1.75 74.22$\pm$4.21
AnatCL-L7 (Destrieux) 68.06$\pm$4.93 70.51$\pm$10.15 63.73$\pm$17.27 60.56$\pm$14.39 55.79$\pm$2.25 59.68$\pm$1.02 56.28$\pm$3.90 81.10$\pm$2.21 63.92$\pm$5.43 77.33$\pm$4.91
AnatCL-G7 (Desikan) 68.16$\pm$5.95 74.06$\pm$8.34 61.97$\pm$15.20 69.33$\pm$19.83 55.60$\pm$2.37 58.04$\pm$5.92 57.15$\pm$6.08 80.79$\pm$1.32 63.46$\pm$3.46 75.69$\pm$3.22
AnatCL-G7 (Destrieux) 66.98$\pm$3.32 72.39$\pm$6.55 59.66$\pm$16.00 67.04$\pm$14.83 54.47$\pm$3.96 60.85$\pm$5.79 55.99$\pm$6.90 79.08$\pm$2.85 64.45$\pm$2.13 75.93$\pm$3.81
OpenBHB provides 7 different anatomical measures, computed with FreeSurfer: Cortical Thickness (CT, mean and std values), Gray Matter Volume (GMV), Surface Area, Integrated Mean, Gaussian Curvature Index, and Intrinsic Curvature Index. The features we select are inspired by the MIND metric (Sebenius et al., 2023). To further motivate the choice of the anatomical features, we conduct a preliminary study by training a simple linear model (Ridge regression) over each feature, with the task of brain age prediction on the OpenBHB data. The results are evaluated in terms of negative mean absolute error and $R^{2}$ score and are reported in Tab. 4. The best results are achieved by GVM and Surface Area, which are morphological features regarding the volume and the total area of each ROI. Thus, we select these two features for our experiments with AnatCL. In addition to GMV and Surface Area measures, we also include CT as it is a widely used measure in related works (Sihag et al., 2024a, b). The final choice is thus composed of three features, which cover different aspects of the morphology and geometry of brain ROIs (e.g., cortical thickness is relevant for conditions such as Alzheimer’s Disease (Phan et al., 2024), and volume and surface measurements can signal brain atrophy, linked to different conditions).
In addition, we also report results of AnatCL employing all seven available measures in Tab. 5. As we can observe, using just the chosen features or all available ones achieves comparable results, with some strongest peak in the reduced version.